# Evaluating machine learning models for predicting pesticides toxicity to honey bees

### v3

@@ 2.3.5 Pretrained neural networks
Those models can be finetuned, using the pretraining model as a weight initialization [69]. However, this approach assumes that there is enough data available to avoid overfitting the model, which is often not the case for highly specific domains in chemistry. Instead, one can use the molecule embeddings from the pretrained model, like a molecular fingerprint, and input it to any conventional ML model. This approach is equivalent to freezing most model weights and training only the classification head. It is also known as frozen embeddings [69], and has shown results competitive with finetuning models [70], while being faster and less prone to overfitting.
Molecule Attention Transformer [65] and its later extension, the Relative Molecule Self-attention Transformer (R-MAT) [71], incorporate many inductive biases specific to molecular chemistry into the GNN graph transformer and their pre-training process. They are based on self-attention between atoms, which is a weighted sum of three components, incorporating molecule-specific inductive biases. In particular, both utilize key-value attention between atom embeddings and pairwise shortest paths between atoms, incorporating both local and global molecule context. Both follow the encoder-only transformer architecture (similar to BERT [72]), interleaving layers of attention, normalization, and MLPs. The learned atom embeddings are combined with the average readout in MAT and with the attention-based pooling in R-MAT. The pretraining procedure for MAT is based only on masked node reconstruction [73], which is a graph-based equivalent of masked language modeling, inspired by BERT. R-MAT uses more challenging masked subgraph reconstruction (similar to GROVER [74]), and multitask regression at the molecule level (inspired by MolBERT [70]).
@@
A notable family of novel molecular models are transformers based solely on SMILES, as exemplified by ChemBERTa [75, 76]. It follows the decoder-only architecture and can be pretrained only on masked language modeling (MLM) on SMILES. Alternatively, it can be pretrained on multitask regression (MTR), predicting many physicochemical properties of molecules computed with RDKit. BERT-like models have relatively simple architectures but result in highly performant embeddings.
@@ 2.3.6 Explainability
We focus on a local, model-agnostic interpretability methods [78], and specifically counterfactual explanations. Those methods interpret a prediction made for a single molecule with an easy visualization of the results. They are independent of the model used, which is very useful as we explore a wide variety of approaches to molecular property prediction. For binary classification, as in ApisTox, counterfactual explanations are very useful, answering the following question: “What is the minimal change required for the input molecule so that the model changes the prediction to the opposite class?”. Algorithms of this type can suggest substituents, functional groups, or fragment removal from the input molecule, so that the model changes the prediction, e.g. from non-toxic to toxic.
@@ 2.4.1 Models implementation
We implement molecular fingerprints using the scikit-fingerprints library [81], using all 2D fingerprints available in the library version 1.12.0. We omit 3D conformer-based fingerprints, since they gave very subpar results in initial experiments. For all fingerprints, we use Random Forest classifier from scikit-learn [82], with 100 trees (default value) and entropy as split criterion. Entropy worked slightly better than the default Gini entropy in our initial experiments, in agreement with previous works [33, 34, 83]. For model regularization, we set the minimal number of samples to split.
@@
GNNs are implemented in PyTorch [85] and PyTorch Geometric [86]. We implement GCN, GraphSAGE, GIN, GAT, and AttentiveFP. Following many architectures in a popular OGB benchmark [87], we extract a standard set of features for atoms and bonds, and use the embedding layer to obtain their initial vectors. Atom features are based on invariants used in ECFP, e.g. atomic number, degree, formal charge. However, we note that only GIN, GAT, and AttentiveFP make use of bond features. We also use Jumping Knowledge (concatenation variant) [62], LayerNorm [63, 64]. For readout, we use sum pooling, proven to be the most expressive [59]. We make an exception for AttentiveFP, which uses its original GRU-based pooling instead. We use the same number of channels in each layer and treat the number of layers and channels as hyperparameters, while also tuning the learning rate and dropout. Networks are trained for 100 epochs, without early stopping.
@@ 2.4.2 Evaluation
For a realistic and challenging train-test split, we use two splits provided with ApisTox: MaxMin split and time split approximation. The test set in MaxMin split is created by selecting maximally diverse molecules, i.e. with the highest sum of pairwise Tanimoto distances between their ECFP4 fingerprints. This way, we require the toxicity classifier to perform well across the whole chemical space of the dataset. Since the maximum diversity picking problem is NP-hard [88], MaxMin heuristic from RDKit is used here [16]. Time split is approximated by PubChem literature data, with the test set consisting of the newest agrochemicals. This requires out-of-distribution generalization capabilities, since new pesticides are typically structurally novel. For both splits, the test set is 20% of the data.
To validate the performance of the model, we measure multiple metrics, appropriate for imbalanced datasets: AUROC, precision, recall, and Matthews correlation coefficient (MCC) [89, 90, 91]. Comparison of models based on multiple metrics is important, as they highlight potential differences in their performance. In particular, MCC is reported to better reflect model performance compared to AUROC [89], and precision and recall focus on false positives and false negatives, respectively.
@@ 3 Data and software availability
We release our code publicly at https://github.com/j-adamczyk/ApisTox_bee_toxicity_ML_prediction. The ApisTox dataset, along with MaxMin and time splits, is available on Zenodo [93] and GitHub [94]. We also include it in our GitHub repository with code for reproducing experimental results.
@@ 4 Results and discussion
Molecular fingerprints are, overall, the most performant and robust solution. Interestingly, quite different fingerprints are the best for both splits. For MaxMin, the best fingerprint is Avalon [95], with 3 substructural fingerprints (Laggner [20], MACCS [21], PubChem [24]) also among the top 5. However, for time split, all best fingerprints are hashed, e.g. ECFP [23], Layered [92], RDKit [92]. This confirms that flexible, more data-driven hashed fingerprints are better suited for out-of-distribution generalization. In contrast, in-domain predictions can be made accurately by extracting common functional groups and other substructural patterns. Since a very large number of fingerprint algorithms are available, understanding such implications can be useful for practical applications.
The performance of GNNs trained from scratch is very underwhelming in both cases. All are unable to outperform baselines, even the simplest atom count in the MaxMin split. This is despite significant computational resources and extensive hyperparameter tuning, even for AttentiveFP, designed specifically for molecular property prediction. However, GNNs are known to require larger amounts of data to train effectively, and by its nature, agrochemistry requires low-data ML. Furthermore, it frequently involves salts, mixtures, and other molecules with disconnected components in their molecular graphs [14, 96, 97]. This is problematic for message-passing models, since they can obtain this information only in the last readout layer, limiting the effectiveness of node embedding learning from their neighborhoods.
@@ 4.1.3 Share of fragmented molecules
Fragmented molecules are defined as those in which two or more independent fragments are present in a single SMILES record. Examples include salts or mixtures, frequently found in agrochemical QSAR [98, 99]. Such compounds can be challenging for some methods, especially those based on graph neural networks (GNNs) or path-based fingerprints like Atom Pair. In such cases, the individual fragments are treated as separate graphs, which prevents the message passing between them. This can impact the performance of the models.
@@ 4.1.4 Unique functional groups
The functional groups defined in the Laggner fingerprint [20] were utilized. It is a collection of 307 fragments, also including those typically indicating toxicity or generally less common in typical drug-like collections.
@@ 4.3 Interpretability - uncertain molecules
In addition to the main dataset and its splits, ApisTox provides a smaller dataset of molecules rejected from the main dataset due to disagreements between experimental measurements in the ECOTOX database. For the main dataset, only molecules with perfectly agreeing $LD_{50}$ values were included, i.e., all under or over $11$ $\mu\text{g / bee}$ (US EPA guidelines threshold [103]). Others are provided as a file with uncertain molecules. Performing predictions on those compounds provides additional evidence, in addition to the literature and experiments, about their toxicity. This is also a very practical application of proposed models, as conflicting measurements can easily arise in ecotoxicology research [104, 105], e.g. due to variations in experimental procedures or outliers in results.
@@
We conclude that the ML model generally agrees with the majority of experimental data, which is expected from a reasonable method. At the same time, it was also able to find more nuanced relations related to real-world regulatory status, which were not obvious from the data. Such applications of predictive modeling may be used as additional evidence for regulatory purposes.
@@ 5 Conclusions
This study provides a comprehensive evaluation of machine learning models to predict pesticide toxicity to honey bees using the ApisTox dataset. Our findings highlight the distinct chemical space occupied by agrochemical compounds compared to medicinal datasets, emphasizing the need for domain-specific predictive models.
@@
Our analysis of the properties and diversity of the datasets revealed significant differences between pesticides and medicinal drug-like compounds, reinforcing the necessity of expanding benchmarks beyond the established medicinal chemistry applications.

### v4

@@ 2.3.5 Pretrained neural networks
Those models can be finetuned, using the pretraining model as a weight initialization [52]. However, this approach assumes that there is enough data available to avoid overfitting the model, which is often not the case for highly specific domains in chemistry. Instead, one can use the molecule embeddings from the pretrained model, like a molecular fingerprint, and input it to any conventional ML model. This approach is equivalent to freezing most model weights and training only the classification head. It is also known as frozen embeddings, and has shown results competitive with finetuning models, while being faster and less prone to overfitting [52].
Molecule Attention Transformer [33] and its later extension, the Relative Molecule Self-attention Transformer (R-MAT) [34], incorporate many inductive biases specific to molecular chemistry into the GNN graph transformer and their pre-training process. They are based on self-attention between atoms, which is a weighted sum of three components, incorporating molecule-specific inductive biases. In particular, both utilize key-value attention between atom embeddings and pairwise shortest paths between atoms, incorporating both local and global molecule context. Both follow the encoder-only transformer architecture, interleaving layers of attention, normalization, and MLPs. The learned atom embeddings are combined with the average readout in MAT and with the attention-based pooling in R-MAT. The pretraining procedure for MAT is based only on masked node reconstruction [33], which is a graph-based equivalent of masked language modeling. R-MAT uses more challenging masked subgraph reconstruction (similar to GROVER [51]), and multitask regression at the molecule level.
@@
A notable family of novel molecular models are transformers based solely on SMILES, as exemplified by ChemBERTa [11]. It follows the decoder-only architecture and can be pretrained only on masked language modeling (MLM) on SMILES. Alternatively, it can be pretrained on multitask regression (MTR), predicting many physicochemical properties of molecules computed with RDKit. BERT-like models have relatively simple architectures but result in highly performant embeddings.
@@ 2.3.6 Explainability
We focus on a local, model-agnostic interpretability methods, and specifically counterfactual explanations. Those methods interpret a prediction made for a single molecule with an easy visualization of the results. They are independent of the model used, which is very useful as we explore a wide variety of approaches to molecular property prediction. For binary classification, as in ApisTox, counterfactual explanations are very useful, answering the following question: “What is the minimal change required for the input molecule so that the model changes the prediction to the opposite class?”. Algorithms of this type can suggest substituents, functional groups, or fragment removal from the input molecule, so that the model changes the prediction, e.g. from non-toxic to toxic.
@@ 2.4.1 Models implementation
We implement molecular fingerprints using the scikit-fingerprints library [3], using all 2D fingerprints available in the library version 1.12.0. We omit 3D conformer-based fingerprints, since they gave very subpar results in initial experiments. For all fingerprints, we use Random Forest classifier from scikit-learn [44], with 100 trees (default value) and entropy as split criterion. Entropy worked slightly better than the default Gini entropy in our initial experiments, in agreement with previous works [1, 2]. For model regularization, we set the minimal number of samples to split.
@@
GNNs are implemented in PyTorch [42] and PyTorch Geometric [18]. We implement GCN, GraphSAGE, GIN, GAT, and AttentiveFP. Following many architectures in a popular OGB benchmark [23], we extract a standard set of features for atoms and bonds, and use the embedding layer to obtain their initial vectors. Atom features are based on invariants used in ECFP, e.g. atomic number, degree, formal charge. However, we note that only GIN, GAT, and AttentiveFP make use of bond features. We also use Jumping Knowledge (concatenation variant) [69] and LayerNorm normalization [33]. For readout, we use sum pooling, proven to be the most expressive [68]. We make an exception for AttentiveFP, which uses its original GRU-based pooling instead. We use the same number of channels in each layer and treat the number of layers and channels as hyperparameters, while also tuning the learning rate and dropout. Networks are trained for 100 epochs, without early stopping.
@@ 2.4.2 Evaluation
For a realistic and challenging train-test split, we use two splits provided with ApisTox: MaxMin split and time split approximation. The test set in MaxMin split is created by selecting maximally diverse molecules, i.e. with the highest sum of pairwise Tanimoto distances between their ECFP4 fingerprints. This way, we require the toxicity classifier to perform well across the whole chemical space of the dataset. Since the maximum diversity picking problem is NP-hard, MaxMin heuristic from RDKit is used here [4]. Time split is approximated by PubChem literature data, with the test set consisting of the newest agrochemicals. This requires out-of-distribution generalization capabilities, since new pesticides are typically structurally novel. For both splits, the test set is 20% of the data.
To validate the performance of the model, we measure multiple metrics, as recommended in ApisTox paper [4]. We select metrics appropriate for imbalanced datasets: AUROC, precision, recall, and Matthews correlation coefficient (MCC). Comparison of models based on multiple metrics is important, as they highlight potential differences in their performance. In particular, MCC is reported to better reflect model performance compared to AUROC, and precision and recall focus on false positives and false negatives, respectively.
@@ 3 Data and software availability
We release our code publicly at https://github.com/j-adamczyk/ApisTox_bee_toxicity_ML_prediction. The ApisTox dataset, along with MaxMin and time splits, is available on Zenodo (https://doi.org/10.5281/zenodo.13350981) and GitHub (https://github.com/j-adamczyk/ApisTox_dataset). We also include it in our GitHub repository with code for reproducing experimental results.
@@ 4 Results and discussion
Molecular fingerprints are, overall, the most performant and robust solution. Interestingly, quite different fingerprints are the best for both splits. For MaxMin, the best fingerprint is Avalon, with 3 substructural fingerprints (Laggner, MACCS, PubChem) also among the top 5. However, for time split, all best fingerprints are hashed, e.g. ECFP, Layered, RDKit. This confirms that flexible, more data-driven hashed fingerprints are better suited for out-of-distribution generalization. In contrast, in-domain predictions can be made accurately by extracting common functional groups and other substructural patterns. Since a very large number of fingerprint algorithms are available, understanding such implications can be useful for practical applications.
The performance of GNNs trained from scratch is very underwhelming in both cases. All are unable to outperform baselines, even the simplest atom count in the MaxMin split. This is despite significant computational resources and extensive hyperparameter tuning, even for AttentiveFP, designed specifically for molecular property prediction. However, GNNs are known to require larger amounts of data to train effectively, and by its nature, agrochemistry requires low-data ML. Furthermore, it frequently involves salts, mixtures, and other molecules with disconnected components in their molecular graphs [4]. This is problematic for message-passing models, since they can obtain this information only in the last readout layer, limiting the effectiveness of node embedding learning from their neighborhoods.
@@ 4.1.3 Share of fragmented molecules
Fragmented molecules are defined as those in which two or more independent fragments are present in a single SMILES record. Examples include salts or mixtures, frequently found in agrochemical QSAR [4]. Such compounds can be challenging for some methods, especially those based on graph neural networks (GNNs) or path-based fingerprints like Atom Pair. In such cases, the individual fragments are treated as separate graphs, which prevents the message passing between them. This can impact the performance of the models.
@@ 4.1.4 Unique functional groups
The functional groups defined in the Laggner fingerprint were utilized. It is a collection of 307 fragments, also including those typically indicating toxicity or generally less common in typical drug-like collections. The full list, including SMARTS definitions, is available at https://github.com/scikit-fingerprints/scikit-fingerprints/blob/master/skfp/fingerprints/data/SMARTS_InteLigand.txt.
@@ 4.3 Interpretability - uncertain molecules
In addition to the main dataset and its splits, ApisTox provides a smaller dataset of molecules rejected from the main dataset due to disagreements between experimental measurements in the ECOTOX database. For the main dataset, only molecules with perfectly agreeing $LD_{50}$ values were included, i.e., all under or over $11$ $\mu\text{g / bee}$ (threshold used in US EPA guidelines). Others are provided as a file with uncertain molecules. Performing predictions on those compounds provides additional evidence, in addition to the literature and experiments, about their toxicity. This is also a very practical application of proposed models, as conflicting measurements can easily arise in ecotoxicology research [4], e.g. due to variations in experimental procedures or outliers in results.
@@
In conclusion, the created ML model generally agrees with the majority of experimental data, which is expected from a reasonable method. At the same time, it was also able to find more nuanced relations related to real-world regulatory status, which were not obvious from the data. Such applications of predictive modeling may be used as additional evidence for regulatory purposes.
@@ 5 Conclusions
This study provides a comprehensive evaluation of machine learning models to predict pesticide toxicity to honey bees using the ApisTox dataset. Findings highlight the distinct chemical space occupied by agrochemical compounds compared to medicinal datasets, emphasizing the need for domain-specific predictive models.
@@
Analysis of the properties and diversity of ApisTox dataset revealed significant differences between pesticides and medicinal drug-like compounds, reinforcing the necessity of expanding benchmarks beyond the established medicinal chemistry applications.